OpenStudy (anonymous):

What is the solution set of |x – 4| + 7 = 4? –4 or 4 –3 or 3 No solution 1 or 7

Parth (parthkohli):

First subtract 7 from both sides. \( \color{Black}{\Rightarrow |x - 4| = -3 }\) We have: \( \color{Black}{\Rightarrow x - 4 = \pm 3 }\)

Parth (parthkohli):

Solve the two equations: \( \color{Black}{\Rightarrow x - 4 = 3 }\) \( \color{Black}{\Rightarrow x - 4 = -3 }\)

OpenStudy (anonymous):

no solution

Parth (parthkohli):

Correct. I was testing you ^_^